OpenStudy (anonymous):

Solve Eleven times x equals fifty-five for x. 5 forty-four 4 fifty-five

OpenStudy (anonymous):

11 times what equal 55 ?

OpenStudy (austinl):

\(11\times x=55\) Do you have any thoughts? How do you cancel a multiplication?

OpenStudy (anonymous):

11*5=55 Answer is 5

OpenStudy (anonymous):

Don't give the answer if they don't even bother responding to those trying to help!! Soooo annoying.

OpenStudy (dobby1):

the answer is 5 because 55/11 is 5 so x =5

OpenStudy (anonymous):

guys i figured it out right after i posted the question

OpenStudy (austinl):

Then close the question.

OpenStudy (anonymous):

well stop looking at it

OpenStudy (anonymous):

Lol ^